OpenStudy (anonymous):

2x - y = 9 3x + 4y = -14 solve for substitution and elimination??

OpenStudy (anonymous):

-2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 2.Solve using substitution- x-y-z=-8 -4x+4y+5z=7 2x+2z=4 5y+6z=5 (-) 9y+6z=9 (=) -4y=-4 y=1 Now that we have a value of y, you can substitute